A capacitor consists of two square metal plates, each measuring 5.00 π‘₯ 10^βˆ’2m on a


side. In between the plates is a sheet of mica measuring 1. 00 π‘₯ 10^βˆ’4m thick.



(a) What is the capacitance of this capacitor? If the charge in one plate is


2. 00 π‘₯ 10^βˆ’8C.




(b) What is the potential difference and




(c) What is the electric field between the plates?



Given: Permittivity of Mica


(πœ– = 4.8 π‘₯ 10^βˆ’11𝐢^2 / 𝑁. π‘š^2)

1
Expert's answer
2022-03-23T07:15:27-0400

(a) What is the capacitance of this capacitor?

"C=\frac{\epsilon A}{d}\\
=\frac{4.8\times10^{-11}\times 5.00\times 10^{-2}}{1.00\times 10^{-4}}=2.40\times 10^{-8}\:\rm F"

(b) Β What is the potential difference between plates

"V=\frac{q}{C}\\
=\frac{2.00\times 10^{-8}}{2.40\times 10^{-8}}=0.833\:\rm V"

(c) What is the electric field between the plates?

"E=\frac{V}{d}\\
=\frac{0.833}{1.00\times 10^{-4}}=8330\:\rm V/m"